The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

           mcnews@astro.uiuc.edu
http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
   http://www.astro.uni-bonn.de/~mcnews/

No. 47 September 14, 2000

Contents

Abstracts of 8 refereed papers  1
Abstracts of 5 non-refereed papers 7
1 job opportunity  10
1 meeting announcement  11

Note from the editors: We apologize that due to time constraints (travels, proposals, and teaching duties) we were unable to send out a call for abstracts before this newsletter was compiled. Calls for abstracts will resume for future editions.

Abstracts of Refereed Papers

HST Spectroscopy of Spot 1

on the Circumstellar Ring of SN 1987A

Eli Michael^{1}, Richard McCray^{1}, C. S. J. Pun^{2}, Peter Garnavich^{3}, Peter Challis^{3}, Robert P. Kirshner^{3}, John Raymond^{3}, Kazimierz Borkowski^{4}, Roger Chevalier^{5}, Alexei V. Filippenko^{6}, Claes Fransson^{7}, Peter Lundqvist^{7}, Nino Panagia^{8}, M. M. Phillips^{9}, George Sonneborn^{2}, Nicholas B. Suntzeff^{10}, Lifan Wang^{11}, and J. Craig Wheeler^{11}

1 JILA, University of Colorado, Boulder, CO 80309-0440, USA 2 Laboratory for Astronomy and Space Physics, Code 681, NASA-GSFC, Greenbelt, MD 20771, USA 3 Harvard-Smithsonian Center for Astrophysics, 60 Garden St, Cambridge, MA 02138, USA 4 Dept. of Physics, North Carolina State University, Raleigh, NC 27695, USA 5 Dept. of Astronomy, University of Virginia, P.O. Box 3818, Charlottesville, VA 22903-0818, USA 6 Dept. of Astronomy, University of California, Berkeley, CA 94720-3411, USA 7 Stockholm Observatory, SE-133 36 Saltsjöbaden, Sweden 8 STScI, 3700 San Martin Drive, Baltimore, MD 21218; on assignment from the Space Science Department of ESA. 9 Las Campanas Observatory, Carnegie Observatories, Casilla 601, La Serena, Chile 10 CTIO, NOAO, Casilla 603, La Serena, Chile 11 Dept. of Astronomy, University of Texas, Austin, TX 78712, USA

We present ultraviolet and optical spectra of the first bright spot (PA = 29^{∘}) on Supernova 1987A’s equatorial circumstellar ring taken with the Space Telescope Imaging Spectrograph. We interpret this spot as the emission produced by radiative shocks that occur where the supernova blast wave strikes an inward protrusion of the ring. The observed line widths and intensity ratios indicate the presence of radiative shocks with velocities ranging from 100 to 250 km s^{−1} entering dense (≥ 10^{4} cm^{−3}) gas. These observations, and future observations of the development of the spectra and line profiles, provide a unique opportunity to study the hydrodynamics of radiative shocks.

Accepted by: Astrophysical Journal Letters

For preprints, contact: michaele@colorado.edu
Also available from the URL: http://xxx.lanl.gov/abs/astro-ph/0008130

The Supergiant Shell LMC2: II. Physical Properties of the 10^{6} K Gas

S.D. Points^{1}, Y.-H. Chu^{1}, S.L. Snowden^{2}, and L. Staveley-Smith^{3}

1 University of Illinois at Urbana-Champaign, USA 2 NASA’s Goddard Space Flight Center, USA 3 Australia Telescope National Facility, Australia

LMC2 has the highest X-ray surface brightness of all know supergiant shells in the Large Magellanic Cloud (LMC). The X-ray emission peaks within the ionized filaments that define the shell boundary, but also extends beyond the southern border of LMC2 as an X-ray bright spur. ROSAT HRI images reveal the X-ray emission from LMC2 and the spur to be truly diffuse, indicating a hot plasma origin. We have obtained ROSAT PSPC and ASCA SIS spectra to study the physical conditions of the hot gas interior to LMC2 and the spur. Raymond-Smith thermal plasma model fits to the X-ray spectra, constrained by HI 21-cm emission-line measurements of the column density, show the plasma temperature of the hot gas interior of LMC2 to be kT = 0.1 - 0.7 keV and of the spur to be kT = 0.1 - 0.5 keV. We have compared the physical conditions of the hot gas interior to LMC2 with those of other supergiant shells, superbubbles, and supernova remnants (SNRs) in the LMC. We find that our derived electron densities for the hot gas inside LMC2 is higher than the value determined for the supergiant shell LMC4, comparable to the value determined for the superbubble N11, and lower than the values determined for the superbubble N44 and a number of SNRs.

Accepted by: The Astrophysical Journal, 2001 Jan 1 issue

For preprints, contact points@astro.uiuc.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0008440

The Antarctic Submillimeter Telescope and Remote Observatory (AST/RO)

Antony A. Stark^{1,7}, John Bally^{2,7}, Simon P. Balm^{1}, T. M. Bania^{3}, Alberto D. Bolatto^{3}, Richard A. Chamberlin^{3,4}, Gregory Engargiola^{5,6}, Maohai Huang^{3}, James G. Ingalls^{3,9},

Karl Jacobs^{8}, James M. Jackson^{3}, Jacob W. Kooi^{9}, Adair P. Lane^{1}, K.-Y. Lo^{5}, Rodney D. Marks^{1}, Christopher L. Martin^{1}, Dennis Mumma^{7}, Roopesh Ojha^{1},

Rudolf Schieder^{8}, Johannes Staguhn^{8,10}, Jürgen Stutzki^{8}, Christopher K. Walker^{11}, Robert W. Wilson^{1,7}, Gregory A. Wright^{7}, Xiaolei Zhang^{1,10}, Peter Zimmermann^{12} and

Rüdiger Zimmermann^{12}

1 Harvard Smithsonian Center for Astrophysics; 60 Garden St.; Cambridge, MA 02138 USA 2 CASA; University of Colorado; Boulder, CO 80309, USA 3 Boston University; 725 Commonwealth Ave.; Boston, MA 02215, USA 4 Caltech Submillimeter Observatory; 111 Nowelo St.; Hilo, HI 96720, USA 5 University of Illinois; 1002 W. Green St.; Urbana, IL 61801, USA 6 Radio Astronomy Lab; University of California; Berkeley, CA, 94720 USA 7 Bell Laboratories; 791 Holmdel-Keyport Rd; Holmdel, NJ 07733, USA 8 I. Physikalisches Institut, Universität zu Köln; Zülpicher Straße 77; D-50937 Köln, Germany 9 California Institute of Technology; Pasadena, CA 91125, USA 10 NASA GSFC; Code 685; Greenbelt, MD 20771, USA 11 Steward Observatory; 933 N. Cherry Ave., University of Arizona, Tucson, AZ 8572, USA 12 Radiometer Physics GmbH; Birkenmaarstraße 10; D-53340 Meckenheim, Germany

AST/RO, a 1.7 m diameter telescope for astronomy and aeronomy studies at wavelengths between 200 and 2000 μm, was installed at the South Pole during the 1994-95 Austral summer. The telescope operates continuously through the Austral winter, and is being used primarily for spectroscopic studies of neutral atomic carbon and carbon monoxide in the interstellar medium of the Milky Way and the Magellanic Clouds. The South Pole environment is unique among observatory sites for unusually low wind speeds, low absolute humidity, and the consistent clarity of the submillimeter sky. Four heterodyne receivers and three acousto-optical spectrometers are installed. Telescope pointing, focus, and calibration methods as well as the unique working environment and logistical requirements of the South Pole are described.

Submitted to: The Astrophysical Journal For preprints, contact aas@cfa.harvard.edu

Available from the URL http://xxx.lanl.gov/abs/astro-ph/0008253

ISO Observations of the Planetary Nebula Lindsay 305 in the Small Magellanic Cloud

A. Kučinskas^{1} ^{2}, V. Vansevičius^{3}, M. Sauvage^{4} and T. Tanabé^{5}

1 National Astronomical Observatory, Tokyo, Japan 2 Institute of Theoretical Physics and Astronomy, Vilnius, Lithuania 3 Institute of Physics, Vilnius, Lithuania 4 CEA/DSM/DAPNIA/Service d’Astrophys. C. E. Saclay, Gif-sur-Yvette Cedex, France 5 Institute of Astronomy, The University of Tokyo, Tokyo, Japan

We present ISO (Infrared Space Observatory) observations of the planetary nebula Lindsay 305 (L 305) in the Small Magellanic Cloud. L 305 is well prominent in the ISOCAM frames at 6.75 and 11.5 μm, although it is under the detection limit at 4.5 μm. The obtained spectral energy distribution shows a strong mid-IR excess, which, depending on the amount of energy radiated at wavelengths longer than 11.5 μm, may be as large as ∼ 1500L_{⊙}. However, since an accurate estimate of the total nebular luminosity is not available to date, the evolutionary status of L 305 can not yet be constrained precisely.

Accepted by: Publications of the Astronomical Society of Japan

For preprints, contact arunaskc@cc.nao.ac.jp

Dust Grain Size Distributions and Extinction in the Milky Way, LMC, and SMC

Joseph C. Weingartner^{1} and B. T. Draine^{2}

1 CITA, 60 St. George Street, University of Toronto, Toronto, ON M5S 3H8, Canada 2 Princeton University Observatory, Peyton Hall, Princeton, NJ 08544, USA

We construct size distributions for carbonaceous and silicate grain populations in different regions of the Milky Way, LMC, and SMC. The size distributions include sufficient very small carbonaceous grains (including polycyclic aromatic hydrocarbon molecules) to account for the observed infrared and microwave emission from the diffuse interstellar medium. Our distributions reproduce the observed extinction of starlight, which varies depending upon the interstellar environment through which the light travels. As shown by Cardelli, Clayton & Mathis in 1989, these variations can be roughly parameterized by the ratio of visual extinction to reddening, R_{V}. We adopt a fairly simple functional form for the size distribution, characterized by several parameters. We tabulate these parameters for various combinations of values for R_{V} and b_{C}, the C abundance in very small grains. We also find size distributions for the line of sight to HD 210121, and for sightlines in the LMC and SMC. For several size distributions, we evaluate the albedo and scattering asymmetry parameter, and present model extinction curves extending beyond the Lyman limit.

Submitted to: The Astrophysical Journal

For preprints, contact weingart@cita.utoronto.ca
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0008146

Halo Lensing or LMC Self Lensing?

Insights from the HST CMD of MACHO Microlensing Source Stars

C. Alcock^{1,2}, R.A. Allsman^{3}, D.R. Alves^{1,4}, T.S. Axelrod^{5}, A.C. Becker^{2,6},

D.P. Bennett^{2,7}, K.H. Cook^{1}, A.J. Drake^{1,5}, K.C. Freeman^{5}, M. Geha^{1,9}, K. Griest^{8}, M.J. Lehner^{10}, S.L. Marshall^{1}, D. Minniti^{1,11}, C.A. Nelson^{1,12}, B.A. Peterson^{6},

P. Popowski^{1}, M.R. Pratt^{6}, P.J. Quinn^{13}, C.W. Stubbs^{2,6}, W. Sutherland^{14},
A.B. Tomaney^{6}, T. Vandehei^{8}, D. Welch^{15}

(The MACHO Collaboration)

1 Lawrence Livermore National Laboratory, Livermore, CA 94550, USA 2 Center for Particle Astrophysics, University of California, Berkeley, CA 9472, USA0 3 Supercomputing Facility, Australian National University, Canberra, ACT 0200, Australia 4 Space Telescope Science Institute, 3700 San Martin Dr., Baltimore, MD 21218, USA 5 Research School of Astronomy and Astrophysics, Canberra, Weston Creek, ACT 2611, Australia 6 Departments of Astronomy and Physics, University of Washington, Seattle, WA 98195, USA 7 Department of Physics, University of Notre Dame, IN 46556, USA 8 Department of Physics, University of California, San Diego, CA 92039, USA 9 Department of Astronomy and Astrophysics, University of California, Santa Cruz 95064 , USA 10 Department of Physics, University of Sheffield, Sheffield S3 7RH, UK 11 Depto. de Astronomia, P. Universidad Catolica, Casilla 104, Santiago 22, Chile 12 Department of Physics, University of California, Berkeley, CA 94720, USA 13 European Southern Observatory, Karl Schwarzchild Str. 2, D-8574 8 Gärching bel München, Germany 14 Department of Physics, University of Oxford, Oxford OX1 3RH, U.K. 15 McMaster University, Hamilton, Ontario Canada L8S 4M1

The nature of Large Magellanic Cloud (LMC) microlensing is determined by the location of the microlensing source stars. If the source stars are in the LMC then LMC microlensing is predominantly halo-lensing, if the source stars are located behind the LMC, then LMC microlensing is dominated by self-lensing. We attempt to distinguish between source stars drawn from the average population of the LMC and source stars drawn from a population behind the LMC by examining the HST colormagnitude diagram (CMD) of microlensing source stars. We present WFPC2 HST photometry of eight MACHO microlensing source stars and the surrounding fields in the LMC. The microlensing source stars are identified by deriving accurate centroids in the ground-based MACHO images using difference image analysis (DIA) and then transforming the DIA coordinates to the HST frame. We consider in detail a model for the background population of source stars based on that presented by Zhao, Graff & Guhathakurta. In this model, the source stars have an additional reddening < E(B − V) >= 0.13 mag and a slightly larger distance modulus < Δμ >∼ 0.3 mag than the average LMC population. We also investigate a series of source star models, varying the relative fraction of source stars drawn from the average and background populations and the displacement of the background population from the LMC. The data suggest that not all of the MACHO microlensing source stars are drawn from a background population.

Submitted to: The Astrophysical Journal For preprints, contact cailin@llnl.gov

Also available from the URL http://arXiv.org/abs/astro-ph/0008282

The So-Called “Bar” in the Large Magellanic Cloud

H.S. Zhao^{1} and N.W. Evans^{2}

1 Sterrewacht Leiden, Postbus 9513, 2300 RA Leiden, The Netherlands 2 Theoretical Physics, 1 Keble Rd, Oxford, UK

We propose that the off-centered “bar” in the Large Magellanic Cloud (LMC) is an unvirialized structure slightly misaligned with, and offset from, the plane of the LMC disk. The small displacement and misalignment are consequences of recent tidal interactions with the SMC and the Galaxy. This proposal, though radical, is consistent with the kinematics of the LMC and near-infrared star count maps from the DENIS and 2MASS surveys. It does not violate any of the observational limits on the depth and structure of the LMC – in particular, the reported 25^{∘} − 50^{∘} inclination range of the LMC and the east-west gradient of distance moduli of standard candles. Contributions to LMC microlensing come mainly from the mutual lensing of stars in the disk and the “bar”. Predicted microlensing optical depth is at levels comparable to observation even without including contributions from MACHOs. Possible observations are suggested to discriminate between our misaligned, offset “bar” model and the conventional picture of an off-centered, planar bar.

Accepted by: The Astrophysical Journal (Letters)

Preprints available from the URL http://arXiv.org/abs/astro-ph/0009155

Discovery of 172-s Pulsations from a Be/X-ray Binary Candidate AX J0051.6−7311 in the SMC with ASCA

Jun Yokogawa,^{1} Ken’ichi Torii,^{2} Kensuke Imanishi,^{1} and Katsuji Koyama^{1}

1 Department of Physics, Graduate School of Science, Kyoto University, Sakyo-ku, Kyoto, 606-8502, Japan 2 National Space Development Agency of Japan, 2-1-1 Sengen, Tsukuba, Ibaraki, 305-8505, Japan

The results from three ASCA observations of AX J0051.6−7311 = RX J0051.9−7311 are reported. Coherent pulsations with a barycentric period of 172.40 ± 0.03 s were discovered in the third observation, with an exceptionally long exposure time of ∼ 177 ks. The X-ray spectrum was found to remain unchanged through these observations, with a photon index of ∼ 0.9. Energy-resolved pulse profiles in the third observation reveal that the pulsations are mostly due to photons with an energy above ∼ 2 keV. Archival data of ROSAT and Einstein indicate that AX J0051.6−7311 exhibits a flux variation having a factor > 20.

Accepted by: PASJ

For preprints, contact jun@cr.scphys.kyoto-u.ac.jp
Also available from the URL http://www-cr.scphys.kyoto-u.ac.jp/member/jun/job/

Abstracts of Non-Refereed Papers

The great circle of the Magellanic Stream - the HI observations

N.V.Bystrova^{1}

1 St.Petersburg Branch of the SAO. 196140 St.Petersburg Pulkovskoe shosse 65, SAO, Russia

The RATAN-600 HI observations were made in the new Zone 2 on the Magellanic Stream’s great circle.This Zone is situated azimuthally opposed to the Zone 1 studied earlier.The coordinates of the Zones are: Zone 1 .RA from 22 to 4 hours, DEC from -40 to +62.5 degrees. Zone 2. RA from 10 to 16 hours , DEC also from -40 to +62.5 deg. The preliminary reduction of observations made in 2000 allowed to determine the character of the HI emission in the Zone 2 in order to compare them with the behaviour of the Warps in the Galactic gaseous disk.

To appear in: Conference proceedings of the JENAM- 2000, May 29 - June 3, 2000, Moscow, Russia : “European Astronomy at the Turn of the Millennium”. For preprints, contact N.V.Bystrova bnv@fsao.spb.su

Also available from the URL http://www.inasan.rssi.ru

Some comparisons of the Galactic HI disk and the Magellanic Stream

N.V.Bystrova ^{1}

1 St.Petersburg Branch of the SAO, 196140 St.Petersburg, Pulkovskoe shosse 65, SAO, Russia

Some details of the Warps in the gaseous disk of our Galaxy are compared with the behaviour of the neutral Hydrogen in the Magellanic Stream. The 21 cm radioline observations made in St.Petersburg and Zelenchucskaja are in accordance.The drift curves received with the RATAN-600 radioteledcope lasted 6 or 7 hours and covered two extended Zones choosen azimuthally opposed on the Magellanic Stream’s great circle.The positions and velocities of the HI signals are described and some preliminary conclusions were made in the communication.

To appear in: Proceedings of the JENAM-2000,May 29- June 3, 2000 Moscow, Russia “European Astronomy at the Turn of the Millennium” For preprints, contact N.V.Bystrova bnv@fsao.spb.su

Also available from the URL http://www.inasan.rssi.ru

What is the behavior of the ISM in the SMC?

S. Stanimirovic^{1}

1 Arecibo Observatory, NAIC/Cornell University, HC3 Box 53995, Arecibo, Puerto Rico 00612, USA

We describe quantitatively the neutral hydrogen (HI) and dust content of the interstellar medium (ISM) in the Small Magellanic Cloud (SMC), using the spatial power spectrum. The velocity modification of the HI density power spectrum is investigated and discussed.

To appear in: Proceedings of the 1st Euroconference on “The Evolution of Galaxies. I. Observational Clues” For preprints, contact sstanimi@naic.edu

Mid-infrared studies of the AGB star populations in Local Group galaxies: the Magellanic Clouds

A. Kučinskas^{1} ^{2}, V. Vansevičius^{3}, M. Sauvage^{4} and T. Tanabé^{5}

1 National Astronomical Observatory, Tokyo, Japan 2 Institute of Theoretical Physics and Astronomy, Vilnius, Lithuania 3 Institute of Physics, Vilnius, Lithuania 4 CEA/DSM/DAPNIA/Service d’Astrophys. C. E. Saclay, Gif-sur-Yvette Cedex, France 5 Institute of Astronomy, The University of Tokyo, Tokyo, Japan

We present a brief summary of the ISO survey of AGB stars in 8 populous intermediate age clusters of the Magellanic Clouds. Totally, more than 100 AGB and RGB stars were observed at 4.5, 6.7 and 12 μm. We introduce a new method to estimate effective temperatures of oxygen-rich AGB stars, and derive precise estimates of M_{bol} and T_{eff} for a large population of AGB objects in both LMC and SMC. The derived stellar parameters are used to construct observed HR diagrams which are employed further to estimate cluster ages using the isochrone fitting. We show that proposed method gives a powerful tool to study AGB star populations within the Local Group galaxies, especially, when backed up with the observations from forthcoming mid-IR astronomical space facilities (IRIS, SIRTF) and large ground-based telescopes (SUBARU, VLT, etc.).

To appear in: Proceedings of ISAS International Workshop on Mid- & Far-Infrared Astronomy and Future Space Missions, April 17-18, 2000, ISAS, Japan

For preprints, contact arunaskc@cc.nao.ac.jp

The Local Group

Eva K. Grebel ^{1,2,3}

1 University of Washington, Department of Astronomy, Box 351580, Seattle, WA 98195-1580, USA 2 Hubble Fellow 3 Max-Planck-Institut für Astronomie, Königstuhl 17, D-69117 Heidelberg, Germany

Local Group galaxies such as the Milky Way, the Magellanic Clouds, and M31 are being used by a number of international collaborations to search for microlensing events. Type and number of detections place constraints on dark matter and the stellar populations within and along the line of sight to these galaxies. In this review I briefly discuss the stellar populations, evolutionary histories, and other properties of different types of Local Group galaxies as well as constraints on the dark matter content of these galaxies. Particular emphasis is placed on the dwarf companions of the spiral galaxies in the Local Group.

To appear in: Proceedings of “Microlensing 2000: A New Era of Microlensing Astrophysics,” Cape Town, South Africa, ASP Conf. Ser., eds. J.W. Menzies and P.D. Sackett

For preprints, contact grebel@mpia-hd.mpg.de
Also available from the URL http://arXiv.org/abs/astro-ph/0008249

Job Opportunities

OPEN POSTDOCTORAL AND Ph.D. POSITIONS

Several Postdoctoral and Ph.D. positions are being offered in an exciting new research group at the University of Goettingen, Germany. The positions will be offered for 3 years in the first instance with a possible extension of a further 2 years.

The group focuses on the kinematic and morphological evolution of field galaxies and their transformation in groups and clusters. The new 10m class telescopes will be used (partly in guaranteed time) to observe the Fundamental Plane and Tully Fischer relations of distant field and cluster galaxies. The spectroscopic picture of galaxy/cluster evolution will be enhanced by NIR, HST and XMM/Chandra observations. N-body/SPH simulations of the evolution of individual galaxies in the dense environment of clusters will be directly interrelated with observations. They will both support the interpretation of observed phenomena, as well as give motivation for new observations.

Applications for postdoctoral positions are invited from experienced observers (ideally with good knowledge of multi-object spectroscopy) and from theoreticians working with N-body/SPH simulations, including star and galaxy formation. The start date can be between November 1st, 2000 and April 1st, 2001. Salary is based on German federal public service scales (BAT IIa, about 100,000 DM per year depending on age and marital status) and includes health insurance, a pension scheme etc. It is anticipated that the appointees will initially work about 60with the rest of the time being used to persue their own reseach. Experience with PC networking would be an advantage. In addition, we seek applications from candidates for Ph.D. positions in the above research areas, whose salary will be half BAT IIa.

Goettingen has guaranteed time at the ESO-VLTs, the Hobby-Eberly Telescope in Texas and SALT. It has access to all ESO, ESA (including HST, XMM, Chandra and SOFIA), Calar Alto and McDonald observatory facilities. A modern PC-Linux network is running including GRAPE hardware, and we have regular access to IBM parallel computers.

Applications for Postdoctoral positions should be supported by a curriculum vitae, publication list, 2 letters of recommendation, and a statement on research interests.

Applications for Ph.D. studentships should include copies of university records, a CV and a letter of recommendation.

The University of Goettingen operates an equal opportunities employment policy.

Please send your application until September 30th, 2000, to:

Dr. Bodo Ziegler Universitaetssternwarte Geismarlandstr. 11 37083 Goettingen Germany

For informal inquiries, please contact me by email: bziegler@uni-sw.gwdg.de

Meeting Announcement

Dwarf Galaxies and their Environment

Date: 23 to 27 January 2001

Place: Physikzentrum, Bad Honnef near Bonn, Germany

URL: http://www.astro.uni-bonn.de/∼webgk/dwarfconf.html
Contact: Klaas S. de Boer, deboer@astro.uni-bonn.de

Scientific Goal: Within the framework of the Gradschool The Magellanic Clouds and other Dwarf Galaxies the conference will focus on research developments in the general field of dwarf galaxies and their interactions with the environment. The program will consist of presentations (talks and posters) by invited speakers, by students from the Gradschool, and by further participants.

Our graduate school, funded by the Deutsche Forschungsgemeinschaft, is being jointly run by the Astronomical Institutes of the Universities of Bonn and Bochum since 1993. This very successful education and research programme will enter its last year of funding and we feel this is the right time to get together for a meeting devoted to the topics that our grad school students have been putting their efforts into during the recent past. With our very positive experience of the last workshop of that kind

at Bad Honnef in January 1998 (see http://www.astro.uni-bonn.de/∼webgk/ws98/cover.html)

we are very much looking forward to meet again at this pleasent and comfortable venue.

SOC: Klaas S. de Boer, Rolf Chini, Ralf-Jürgen Dettmar, Uli Klein, Ulrich Mebold, Rheinhard Schlickeiser, Wilhelm Seggewiss

Proceedings: The proceedings of the meeting will be published.

Registration: Thechnical details of the meeting and a registration form can be found on the homepage of the meeting (URL see above). Participants will be selected on a first-come, first-served basis and on how well their proposed contribution fits to the topic of the meeting. The registration deadline is 20 October 2000. Due to space and logistic limitations the number of attendees will be limited to ∼ 75 people.